please help me with this

THANK YOU!!! 1. You are trying to make a homemade electrolyte energy drink similar to Gatorade. Gatorade contains 270. mg of sodium chlori
Answer #1

Let us assume volume of 5% NaCl soln,volume of sucrose soln,volume of extra water added be VNa,Vs,Vw respectively.

Given:

· Density of 5% NaCl soln=1.034g/mL.

· Density of Gatorade =1.030g/mL

· Density of water added=1.034g/mL

· Density of water=1g/mL

· Amount of NaCl in Gatorade =0.270g in 591mL.

· Amount of sugar in Gatorade=34 g in 591mL

· Total volume of solution =240mL=VNa+Vs+Vw

concentration of NaCl in Gatorade, CNaCl in Gatorade =.270/591 g/mL

concentration of sucrose in Gatorade ,CSucrose in Gatorade=34/591 g/mL

Mass of Nacl soln added to make Gatorade=1.034*VNa.

Mass of sucrose soln added to make Gatorade=1.038*Vs.

Mass of sucrose in Gatorade =0.1*mass of sucrose soln.

                                                                =0.1*1.038*Vs

Mass of NaCl in 5 %solution =0.05*mass of nacl solution

                                                              =0.05*1.034*VNa            -(1)

Amount of Nacl in 240 ml Gatorade =concentration of NaCl in Gatorade*volume of solution.

=.270/591*240                 -(2)

Amount ofsucrose in 240 ml Gatorade =concentration of sucrose in Gatorade*volume of solution.

Equating eq(1) and (2)  

=0.05*1.034*VNa=.270/591*240              

VNa=2.12078 mL

Similarly for sucrose

0.1*1.038*Vs=34/591*240

Vs=133.0164 mL.

Also density of solution is =mass of solution/volume of solution.

1.034 VNa+1.038&Vs+Vw VNa+Vs+Vw = 1.03

=1.034*VNa+1.038*Vs+Vw=1.03*240

1.034*VNa+1.038*Vs+Vw=247.2

Vw=106.936mL